OpenStudy (anonymous):

what number must be added in order to make x^2+8x=34 a complete square?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

16

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

always take the middle number divide it by 2 and square it u will get the answer then next step juz add both sides by that number can already

OpenStudy (anonymous):

in this case is 8 divided by 2 =4 then u square the 4 u will get 16

OpenStudy (anonymous):

so a perfect square would be (x+4)^2 and give us the 8x in the middle, to complete the square we have to have -16 on the other side, so we have to subtract 50 (34+16) from the equation.

OpenStudy (anonymous):

then we get x^2+4x=34-50 --> x^2+4x=-16 --> x^2+4x+16=0 --> (x+4)^2=0.

OpenStudy (anonymous):

hoh given equation is x^2+8x-34=0 x^2+2(4)(x)-34=0 adding 50 makes it x^2+8x+16=0 a perfect square (x+2)^2

OpenStudy (anonymous):

you mean (x+4)^2